A body moving with uniform acceleration describes 12 m in 3rd second of its motion and 20 m in 5th second find the velocity after 10 seconds

Grade:NULL

2 Answers

Himanshu Pareek
4 years ago
Displacement in nth second is given as
Sn = u + (a/2)*(2n – 1)
so from the given data
12 = u + 5a/2 eq 1
20 = u + 9a/2 eq 2
by eq 1 and 2
a= 4 m/s^2 and u = 2 m/s
velocity after 10 sec, v = u + a*t
v = 2 + 4*10
v = 42 m/s
Regards,
Himanshu Pareek
Khimraj
4 years ago
As the body is in unform acceleration then the body
velocity per second
4m. 1
8m. 2
12m. 3 
16m. 4
20m. 5
it will get increased by 4m per second
then (4 metres x 10 second)
= 40 m 
therefore the velocity of body is 40m after 10 seconds